OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

do you need both x and y and steps

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so i woul use substion y = -2x + 19 y = x + 7 x(- 2x + 19) + 7 the reason i did this because techonly there is a one there y=-2x+19+7 so combine the 19 and 7 2x+26 So i need to separate the number with a equal sign 2x=26 x=13 now i place this in the equation y=-2(13) +19 y=-26+19 y=-17 (13,-17) i pretty sure

OpenStudy (anonymous):

nah no rig bu thanks for trying